Solve in radians(must work with left side)
1-2sin^2*x=cos^2*x

Answer by stanbon(48558)

Solve in radians(must work with left side)
1-2sin^2*x=cos(2x)
------------------
since sin^2 + cos^2 = 1
Substitute for "1" to get:
(sin^2 + cos^2) - 2sin^2 = cos(2x)
cos^2x - sin^2x = cos(2x)
cos(2x) = cos(2x)
==========================
Cheers,
Stan H.